Suppose we have 2.0 g of ethyl alcohol at -50 degree celsius and we remove heat to it continuously. let us also assume that no heat is lost to the surroundings. How much heat is to be removed to convert this from liquid to solid at -120 degree celsius.


1
Expert's answer
2022-04-20T09:50:03-0400

Moles = 2.0 / 46 = 0.043 mol

From -50 to -114C (freezing point): 118 J/mol*K x 0.043 mol x (114 – 50) = 328 J

At -114C: 4.9 kJ/mol * 0.043 mol = 210.7 J  

From -114 to -120: 118 J/mol*K x 0.043 mol x (120 – 114) = 30.4 J

Together: 328 + 210.7 + 30.4 = 569.1 J